687vhlovdjfl:

Write the following linear equation in function notation. y = 7x - 2

687vhlovdjfl:

starbucksgirl13:

Function Notation and Linear Functions - Algebra | Socraticsocratic.org › algebra › function-notation-and-linear-fu... How do you write linear equations in function notation? First, convert ... y=−3. y=3x−5. y=2x2−3x+1. (1,2),(2,2),(3,2),(4,2). The following are NOT functions: x=3.

Shadow:

y = f(x)

Shadow:

So if I see y = 3x + 5 I can change it to: f(x) = 3x + 5

687vhlovdjfl:

so f(x) = 7x - 2

Shadow:

Yes

687vhlovdjfl:

thank you

Shadow:

You're welcome